# Shuchang Zhang, Hui Zhang, Hongxia Wang

### v4

@@ Abstract
Recently, the degenerate preconditioned proximal point (PPP) method provides a unified and flexible framework for designing and analyzing operator-splitting algorithms such as Douglas-Rachford (DR). However, the degenerate PPP method exhibits weak convergence in the infinite-dimensional Hilbert space and lacks accelerated variants. To address these issues, we propose a Halpern-type PPP (HPPP) algorithm, which leverages the strong convergence and acceleration properties of Halpern’s iteration method. Moreover, we propose a novel algorithm for image restoration by combining HPPP with denoiser priors such as Plug-and-Play (PnP) prior, which can be viewed as an accelerated PnP method. Finally, numerical experiments including several toy examples and image restoration validate the effectiveness of our proposed algorithms.
@@ 1 Introduction
where $f:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ and $g:\mathcal{Y}\to\mathbb{R}\cup\{+\infty\}$ are convex, lower semicontinuous functions, $\mathbf{K}:\mathcal{X}\to\mathcal{Y}$ is a bounded linear operator, and $\lambda>0$ is a balance parameter. Both $\mathcal{X}$ and $\mathcal{Y}$ are real Hilbert spaces. The first term $f$ represents the data fidelity, while the second term $g$ serves as a regularization such as TV [48] to mitigate the ill-posedness of IR problems.
@@
where $\partial f(\mathbf{x})$ and $\partial g(\mathbf{x})$ are the subdifferentials of $f$ and $g$ at $\mathbf{x}$, respectively [5, Chapter 3]. Following [27, 9], by introducing an auxiliary variable $\mathbf{y}\in\partial g(\mathbf{K}\mathbf{x})$, we reformulate (2) as:
@@
where $\mathcal{A}=\begin{pmatrix}\lambda\partial f&\mathbf{K}^{*}\\ -\mathbf{K}&(\partial g)^{-1}\end{pmatrix}$, $\mathbf{u}=(\mathbf{x},\mathbf{y})$, and $\mathcal{H}=\mathcal{X}\times\mathcal{Y}$. The problem (3) is common in modern optimization and variational analysis [2, 3]. When $\mathcal{A}$ is maximal monotone, the resolvent $J_{\mathcal{A}}=(I+\mathcal{A})^{-1}$ is nonexpansive with a full domain, as established by the Minty surjectivity theorem [39]. The proximal point iteration $\mathbf{u}^{k+1}=(I+\mathcal{A})^{-1}\mathbf{u}^{k}$ is used to solve (3) and converges weakly [46]. However, it is difficult to compute the operator $(I+\mathcal{A})^{-1}$, and hence splitting methods have been developed to address this issue. The well-known Douglas-Rachford splitting (DRS) [23] decomposes $\mathcal{A}$ into the sum of two maximal monotone operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, for which $J_{\mathcal{A}_{1}}$ and $J_{\mathcal{A}_{2}}$ are easier to obtain. Another way to solve (1) is to transform it into the saddle-point problem [16, 43], i.e.,
@@
where $g^{*}:\mathcal{K}\to\mathbb{R}\cup\{+\infty\}$ is the conjugate of $g$. In this direction, many primal-dual methods, including the well-known Chambolle-Pock (CP) splitting method [16, 43] and the Primal-Dual Hybrid Gradient (PDHG) method [27], are designed and studied extensively. In particular, He and Yuan [27] analyzed the PDHG method from a PPP standpoint with a positive definite preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$. Recently, Bredies et al. [11, 12, 9] developed a unified degenerate PPP algorithmic framework with a positive semi-definite $\mathcal{M}$. By choosing appropriate preconditioners, this framework could cover the DR and CP splitting algorithms [9].
@@
It is crucial to study algorithms that exhibit both strong convergence and acceleration [30, 8, 34, 3, 52, 17, 38], such as strongly convergent proximal point methods [30] and forward-reflected-backward splitting algorithms [34]. For example, Bauschke et al. demonstrated strong convergence of degenerate PPP under the special case where $\mathcal{A}$ is linear [3]. In such cases, the weak limit of the PPP sequence corresponds to the $\mathcal{M}$-projection of the initial point. Moreover, Sun et al. [52] proposed an accelerated preconditioned Alternating Direction Method of Multipliers (pADMM) by leveraging the degenerate PPP method [9] and the fast Krasnosel’skiĭ-Mann (KM) iteration [6]. Similarly, Chen et al. introduced an accelerated HPR-LP solver, which implements a Halpern Peaceman-Rachford method enhanced with semi-proximal terms for efficiently solving linear programming (LP) problems [17], and the restarted Halpern PDHG (rHPDHG) achieves an accelerated linear convergence rate [38]. A shortcoming of forward-backward and DRS algorithms employing KM iteration is that their iterates converge only in the weak topology [8]. Consequently, the degenerate PPP method based on KM iteration typically exhibits weak convergence in Hilbert spaces. Meanwhile, it is unclear whether the degenerate PPP method can be accelerated via simple modification.
@@
The classic Halpern iteration [26] offers the advantage of strong convergence over the KM iteration in infinite-dimensional Hilbert spaces, with the limit identified as the metric projection of the anchor onto the fixed point set [29]. Then the Halpern iteration is also known as an implicitly regularized method [22]. Due to this implicit regularity, the degenerate PPP method incorporating Halpern iteration can obtain a unique solution, yielding stable recovery results for IR problems. Achieving stable reconstruction for ill-posed inverse problems is important [24]. Moreover, beyond the implicit regularity, Halpern iteration can also accelerate convergence rate in terms of operator residual norm [55], which is widely utilized in machine learning [22, 41, 29]. Notably, the work [41] has demonstrated that PDHG with restarted Halpern iteration achieves a faster convergence rate for function values in CT image reconstruction. Based on the Halpern iteration, the Halpern-type preconditioned proximal point algorithm called HPPP (5) is thus proposed to overcome the two limitations of the degenerate PPP method,
@@
where $\mathbf{a},\mathbf{u}^{0}\in\mathcal{H}$ are the anchor point and the initial point, respectively, and $\{\mu_{k}\}_{k\in\mathbb{N}}$ is a sequence in $[0,1]$ such that $\sum_{k\in\mathbb{N}}\mu_{k}=+\infty,\lim_{k\to\infty}\mu_{k}=0$.
PnP (Plug-and-Play) methods that combine splitting algorithms with denoiser priors have been widely applied in practical problems [56, 51, 1, 35, 59] and have achieved state-of-the-art performance in inverse imaging tasks [65, 33, 53]. Buzzard et al. provided a Consensus Equilibrium interpretation on denoiser priors [14]. Romano et al. introduced RED (Regularization by Denoising) [47], whose gradient exactly corresponds to the denoising residual, thereby yielding a clear objective function [20] that can be exploited in first-order optimization. Powerful denoisers (e.g., DnCNN [66]) typically do not meet the conditions of RED [45]. To address this limitation, Reehorst and Schniter introduced the score-matching by denoising (SMD) perspective to interpret RED [45]. Based on the fixed-point projection, Cohen et al. proposed the RED-PRO model [20] as a bridge between RED and PnP, while the hybrid steepest descent (HSD) method [63, 64] was employed to solve the resulting model. Meanwhile, Ryu et al. proved the convergence of PnP-FBS and PnP-ADMM using the Banach contraction principle under the assumption that the data term $f$ is strongly convex and that the residual is nonexpansive [49]. However, enforcing strong convexity on $f$ precludes many IR tasks [33]. Furthermore, Cohen et al. parameterized denoisers via the gradients of smooth potential functions that satisfy a symmetric Jacobian property [19], and Hurault et al. proposed the gradient step (GS) denoiser [32, 33], which can be interpreted as a proximal operator of an implicit regularizer [33]. Although the convergence of PnP-ADMM is commonly analyzed from the perspective of Douglas-Rachford splitting (DRS) based on their equivalence [49, 33], theoretically achieving fast fixed-point residual decay in PnP-DRS remains unknown.
Based on HPPP and denoiser priors,  [^1] Since the RED gradient exactly equals the residual, we adopt the GraRED notation—derived from Moreau decomposition—to represent the proximal operator of the implicit regularizer associated with the residual.we propose the Gradient Regularization by Denoising via HPPP called GraRED-HP3 for IR problems. The main contributions are as follows:
@@
    Theoretical Contributions. The sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP with the positive semi-definite preconditioner $\mathcal{M}$ converges strongly to a unique solution $\mathbf{u}^{*}=\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\|\mathbf{u}-\mathbf{a}\|_{\mathcal{M}}^{2}$, as stated in Theorem 3.1 and Proposition 3.3, while the original PPP method can only converge weakly to some uncertain solution. Moreover, compared with the asymptotic regularity result of PPP [9, Lemma 2.8] (i.e., $\lim_{k\to\infty}\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|_{\mathcal{M}}=0$), we establish a convergence rate of $\mathcal{O}(1/k)$ for both $\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|$ and $\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|_{\mathcal{M}}$ (see Propositions 3.10 and 3.12).
@@
    Algorithmic Development. We integrate HPPP with denoiser priors to propose the GraRED-HP3 algorithm for IR problems and shed new theoretical insights on denoiser priors from a PPP standpoint. Noting that the PnP-ADMM algorithm can be reformulated as an equivalent DRS form [49, 33], and given that the worst-case convergence rate for DRS is $\mathcal{O}(1/\sqrt{k})$ [28], GraRED-HP3 improved a sublinear convergence rate (as per Proposition 3.12) and thereby serves as an accelerated PnP method.
@@
    Experimental Validation. We numerically validate the advantages of HPPP, including its implicit regularity, acceleration, and efficiency with several toy examples. In addition, we demonstrate the state-of-the-art performance of GraRED-HP3 and its restarted variant with advanced denoisers in IR tasks.
@@ 2.1 Preconditioned proximal point
Bredies et al. [9] introduced the degenerate PPP framework with the positive semi-definite preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$. The proper preconditioner $\mathcal{M}$ can make $\mathcal{A}+\mathcal{M}$ have a lower triangular structure, which conveniently calculates the inverse $(\mathcal{A}+\mathcal{M})^{-1}$.
@@
An admissible preconditioner for the operator $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ is a bouned, linear, self-adjoint, and positive semi-definite operator $\mathcal{M}:\mathcal{H}\to\mathcal{H}$ such that
@@
Therefore, the PPP iteration is written into
@@
If $\mathcal{M}=I$, then $\mathcal{T}$ is firmly nonexpansive (FNE)operator [2, section 23] and (6) becomes the standard proximal point iteration. If $\mathcal{M}$ is positive semi-definite, the operator $\mathcal{T}$ is related with the degenerate $\mathcal{M}$-firmly nonexpansive ($\mathcal{M}$-FNE) operator [11, 15, 62, 9], which is associated with seminorm $\left\|\mathbf{u}\right\|_{\mathcal{M}}=\sqrt{\langle\mathcal{M}\mathbf{u},\mathbf{u}\rangle}$ and semi inner-product $\langle\mathbf{u},\mathbf{v}\rangle_{\mathcal{M}}=\langle\mathcal{M}\mathbf{u},\mathbf{v}\rangle$. The following notion extends monotone characteristics, i.e., $\mathcal{M}$-monotonicity.
@@
Let $\mathcal{M}:\mathcal{H}\to\mathcal{H}$ be a bouned linear positive semi-definite operator, then $\mathcal{B}:\mathcal{H}\to 2^{\mathcal{H}}$ is $\mathcal{M}$-monotone if we have
$\langle\mathbf{v}-\mathbf{v}^{\prime},\mathbf{u}-\mathbf{u}^{\prime}\rangle_{\mathcal{M}}\geq 0,\forall(\mathbf{u},\mathbf{v}),(\mathbf{u}^{\prime},\mathbf{v}^{\prime})\in\mathcal{B}.$
The following lemma demonstrates that the firmly non-expansiveness notation can be generalized to the degenerate case [11, 15]. If $\mathcal{M}$ is an admissible preconditioner and $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone, then the operator $\mathcal{T}$ is $\mathcal{M}$-FNE.
@@
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and $\mathcal{M}$ an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone. Then $\mathcal{T}$ is $\mathcal{M}$-FNE, i.e.,
@@
By the KM iteration, the degenerate PPP method is written into
@@
where $\lambda_{k}$ is a sequence in $[0,2]$ such that $\sum_{k}\lambda_{k}(2-\lambda_{k})=+\infty$.
2.2 PnP and RED

### v5

@@ Abstract
Recently, the degenerate preconditioned proximal point (PPP) method provided a unified and flexible framework for designing and analyzing operator-splitting algorithms, such as the Douglas-Rachford (DR) splitting. However, the degenerate PPP method exhibits weak convergence in the infinite-dimensional Hilbert space and lacks accelerated variants. To address these issues, we propose a Halpern-type PPP (HPPP) algorithm, which leverages the strong convergence and acceleration properties of Halpern’s iteration method. Moreover, we propose a novel algorithm for image restoration by combining HPPP with denoiser priors, such as the Plug-and-Play (PnP) prior, which can be viewed as an accelerated PnP method. Finally, numerical experiments, including several toy examples and image restoration, validate the effectiveness of our proposed algorithms.
@@ 1 Introduction
where $f:\mathcal{X}\to\mathbb{R}\cup\{+\infty\}$ and $g:\mathcal{Y}\to\mathbb{R}\cup\{+\infty\}$ are convex, lower semicontinuous functions, $\mathbf{K}:\mathcal{X}\to\mathcal{Y}$ is a bounded linear operator, and $\lambda>0$ is a balance parameter. Both $\mathcal{X}$ and $\mathcal{Y}$ are real Hilbert spaces. The first term $f$ represents the data fidelity, while the second term $g$ serves as a regularization, such as TV (total variation) [48], to mitigate the ill-posedness of IR problems.
@@
where $\partial f(\mathbf{x})$ and $\partial g(\mathbf{x})$ are the subdifferentials of $f$ and $g$ at $\mathbf{x}$, respectively [5, Chapter 3]. Following [27, 9], by introducing an auxiliary variable $\mathbf{y}\in\partial g(\mathbf{K}\mathbf{x})$, we reformulate (2) as
@@
where $\mathcal{A}=\begin{pmatrix}\lambda\partial f&\mathbf{K}^{*}\\ -\mathbf{K}&(\partial g)^{-1}\end{pmatrix}$, $\mathbf{u}=(\mathbf{x},\mathbf{y})$, and $\mathcal{H}=\mathcal{X}\times\mathcal{Y}$. The problem (3) is common in modern optimization and variational analysis [2, 3]. When $\mathcal{A}$ is maximal monotone, the resolvent $J_{\mathcal{A}}=(I+\mathcal{A})^{-1}$ is nonexpansive with a full domain, as established by the Minty surjectivity theorem [39]. The proximal point iteration $\mathbf{u}^{k+1}=(I+\mathcal{A})^{-1}\mathbf{u}^{k}$ is used to solve (3) and converges weakly [46]. However, it is difficult to compute the operator $(I+\mathcal{A})^{-1}$, and hence splitting methods have been developed to address this issue. The well-known Douglas-Rachford splitting (DRS) [23] decomposes $\mathcal{A}$ into the sum of two maximal monotone operators $\mathcal{A}_{1}$ and $\mathcal{A}_{2}$, for which $J_{\mathcal{A}_{1}}$ and $J_{\mathcal{A}_{2}}$ are easier to obtain. Another way to solve (1) is to transform it into a saddle-point problem [16, 43], i.e.,
@@
where $g^{*}:\mathcal{Y}\to\mathbb{R}\cup\{+\infty\}$ is the conjugate of $g$. In this direction, many primal-dual methods, including the well-known Chambolle-Pock (CP) primal-dual method [16, 43] and the primal-dual hybrid gradient (PDHG) method [27], are designed and studied extensively. In particular, He and Yuan [27] analyzed the PDHG method from a PPP standpoint with a positive definite preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$. Recently, Bredies et al. [11, 12, 9] developed a unified degenerate PPP algorithmic framework with a positive semidefinite $\mathcal{M}$. By choosing appropriate preconditioners, this framework could cover the DR and CP algorithms [9].
@@
It is crucial to study algorithms that exhibit both strong convergence and acceleration [30, 8, 34, 3, 52, 17, 38], such as strongly convergent proximal point methods [30] and forward-reflected-backward splitting algorithms [34]. For example, Bauschke et al. demonstrated strong convergence of degenerate PPP under the special case where $\mathcal{A}$ is linear [3]. In such cases, the weak limit of the PPP sequence corresponds to the $\mathcal{M}$-projection of the initial point. Moreover, Sun et al. [52] proposed an accelerated preconditioned alternating direction method of multipliers (pADMM) by leveraging the degenerate PPP method [9] and the fast Krasnosel’skiĭ-Mann (KM) iteration [6]. Similarly, Chen et al. introduced an accelerated HPR-LP solver, which implements a Halpern Peaceman-Rachford method enhanced with semiproximal terms for efficiently solving linear programming (LP) problems [17], and the restarted Halpern PDHG (rHPDHG) achieves an accelerated linear convergence rate [38]. A shortcoming of forward-backward and DRS algorithms employing KM iteration is that their iterates converge only in the weak topology [8]. Consequently, the degenerate PPP method based on KM iteration typically exhibits weak convergence in Hilbert spaces. Meanwhile, it is unclear whether the degenerate PPP method can be accelerated via simple modification.
@@
The classic Halpern iteration [26] offers the advantage of strong convergence over the KM iteration in infinite-dimensional Hilbert spaces, with the limit identified as the metric projection of the anchor onto the fixed point set [29]. Then the Halpern iteration is also known as an implicitly regularized method [22]. Due to this implicit regularity, the degenerate PPP method incorporating Halpern iteration can obtain a unique solution, yielding stable recovery results for IR problems. Achieving stable reconstruction for ill-posed inverse problems is important [24]. Moreover, beyond the implicit regularity, Halpern iteration can also accelerate convergence rate in terms of operator residual norm [55], which is widely utilized in machine learning [22, 41, 29]. Notably, the work [41] has demonstrated that PDHG with restarted Halpern iteration achieves a faster convergence rate for function values in CT image reconstruction. Based on the Halpern iteration, the following Halpern-type preconditioned proximal point algorithm (called HPPP) (5) is thus proposed to overcome the two limitations of the degenerate PPP method:
@@
where $\mathcal{T}=(\mathcal{A}+\mathcal{M})^{-1}\mathcal{M}$, and $\mathcal{A},\mathcal{M}$ are detailed in subsection 3.3, and $\mathbf{a},\mathbf{u}^{0}\in\mathcal{H}$ are the anchor point and the initial point, respectively, and $\{\mu_{k}\}_{k\in\mathbb{N}}$ is a sequence in $[0,1]$ such that $\sum_{k\in\mathbb{N}}\mu_{k}=+\infty,\lim_{k\to\infty}\mu_{k}=0$.
PnP (Plug-and-Play) methods that combine splitting algorithms with denoiser priors have been widely applied in practical problems [56, 51, 1, 35, 59] and have achieved state-of-the-art performance in inverse imaging tasks [65, 33, 53]. Buzzard et al. provided a consensus equilibrium interpretation on denoiser priors [14]. Romano , Elad, and Milanfar introduced RED (regularization by denoising) [47], whose gradient exactly corresponds to the denoising residual, thereby yielding a clear objective function [20] that can be exploited in first-order optimization. Powerful denoisers such as denoising convolutional neural network (DnCNN) [66] typically do not meet the conditions of RED [45]. To address this limitation, Reehorst and Schniter introduced the score-matching by denoising (SMD) perspective to interpret RED [45]. Based on the fixed-point projection, Cohen et al. proposed the RED-PRO model [20] as a bridge between RED and PnP, while the hybrid steepest descent (HSD) method [63, 64] was employed to solve the resulting model. Meanwhile, Ryu et al. proved the convergence of PnP-FBS (forward-backward splitting) and PnP-ADMM using the Banach contraction principle under the assumption that the data term $f$ is strongly convex and that the residual is nonexpansive [49]. However, enforcing strong convexity on $f$ precludes many IR tasks [33]. Furthermore, Cohen et al. parameterized denoisers via the gradients of smooth potential functions that satisfy a symmetric Jacobian property [19], and Hurault, Leclaire, and Papadakis developed the gradient step (GS) denoiser [32, 33], which can be interpreted as a proximal operator of an implicit regularizer [33]. Although the convergence of PnP-ADMM is commonly analyzed from the perspective of Douglas-Rachford splitting (DRS) based on the equivalence between ADMM and DRS [49, 33], theoretically achieving fast fixed-point residual decay in PnP-DRS remains unknown.
Based on HPPP and PnP priors, we propose the gradient regularization by denoising via HPPP called GraRED-HP3 (Algorithm 3111Since the RED gradient exactly equals the residual, we adopt the GraRED notation—derived from Moreau decomposition—to represent the proximal operator of the implicit regularizer associated with the residual.. Our main contributions are as follows:
@@
    Theoretical contributions. The sequence $\{\mathbf{u}^{k}\}_{k\in\mathbb{N}}$ generated by HPPP with the positive semidefinite preconditioner $\mathcal{M}$ converges strongly to a unique solution $\mathbf{u}^{*}=\arg\min_{\mathbf{u}\in\mathrm{Fix}(\mathcal{T})}\|\mathbf{u}-\mathbf{a}\|_{\mathcal{M}}^{2}$, as stated in Theorem 3.1 and Proposition 3.3, while the original PPP method can only converge weakly to some uncertain solution. Moreover, compared with the asymptotic regularity result of PPP [9, Lemma 2.8] (i.e., $\lim_{k\to\infty}\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|_{\mathcal{M}}=0$), we establish a convergence rate of $\mathcal{O}(1/k)$ for both $\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|$ and $\|\mathcal{T}\mathbf{u}^{k}-\mathbf{u}^{k}\|_{\mathcal{M}}$ (see Propositions 3.10 and 3.12).
@@
    Algorithmic development. We integrate HPPP with denoiser priors to propose the GraRED-HP3 algorithm for IR problems and shed new theoretical insights on denoiser priors from a PPP standpoint. Noting that the PnP-ADMM algorithm can be reformulated in an equivalent DRS form [49, 33], and given that the worst-case convergence rate for DRS is $\mathcal{O}(1/\sqrt{k})$ [28], GraRED-HP3 improved a sublinear convergence rate (as per Proposition 3.12) and thereby serves as an accelerated PnP method.
@@
    Experimental validation. We numerically validate the advantages of HPPP, including its implicit regularity, acceleration, and efficiency, with several toy examples. In addition, we demonstrate the state-of-the-art performance of GraRED-HP3 and its restarted variant with advanced denoisers in IR tasks.
@@ 2.1 Preconditioned proximal point
Bredies et al. [9] introduced the degenerate PPP framework with the positive semidefinite preconditioner $\mathcal{M}:\mathcal{H}\to\mathcal{H}$. The proper preconditioner $\mathcal{M}$ can make $\mathcal{A}+\mathcal{M}$ have a lower triangular structure, which conveniently calculates the inverse $(\mathcal{A}+\mathcal{M})^{-1}$.
@@
An admissible preconditioner for the operator $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ is a bounded, linear, self-adjoint, and positive semidefinite operator $\mathcal{M}:\mathcal{H}\to\mathcal{H}$ such that
@@
Therefore, the PPP iteration is written as
@@
If $\mathcal{M}=I$, then $\mathcal{T}$ is a firmly nonexpansive (FNE) operator [2, Proposition 23.8] and (6) becomes the standard proximal point iteration. If $\mathcal{M}$ is positive semidefinite, the operator $\mathcal{T}$ is related with the degenerate $\mathcal{M}$-firmly nonexpansive ($\mathcal{M}$-FNE) operator [11, 15, 62, 9], which is associated with seminorm $\left\|\mathbf{u}\right\|_{\mathcal{M}}=\sqrt{\langle\mathcal{M}\mathbf{u},\mathbf{u}\rangle}$ and semi inner-product $\langle\mathbf{u},\mathbf{v}\rangle_{\mathcal{M}}=\langle\mathcal{M}\mathbf{u},\mathbf{v}\rangle$. The following notion extends monotone characteristics, i.e., $\mathcal{M}$-monotonicity.
@@
Let $\mathcal{M}:\mathcal{H}\to\mathcal{H}$ be a bounded linear positive semidefinite operator; then $\mathcal{B}:\mathcal{H}\to 2^{\mathcal{H}}$ is $\mathcal{M}$-monotone if we have
$\langle\mathbf{v}-\mathbf{v}^{\prime},\mathbf{u}-\mathbf{u}^{\prime}\rangle_{\mathcal{M}}\geq 0\quad\forall(\mathbf{u},\mathbf{v}),(\mathbf{u}^{\prime},\mathbf{v}^{\prime})\in\mathcal{B}.$
The following lemma demonstrates that the firmly nonexpansive notation can be generalized to the degenerate case [11, 15]. If $\mathcal{M}$ is an admissible preconditioner and $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone, then the operator $\mathcal{T}$ is $\mathcal{M}$-FNE.
@@
Let $\mathcal{A}:\mathcal{H}\to 2^{\mathcal{H}}$ be an operator with $\mathrm{zer}\mathcal{A}\neq\emptyset$, and let $\mathcal{M}$ be an admissible preconditioner such that $\mathcal{M}^{-1}\mathcal{A}$ is $\mathcal{M}$-monotone. Then $\mathcal{T}$ is $\mathcal{M}$-FNE, i.e.,
@@
By the KM iteration, the degenerate PPP method is written as
@@
where $\lambda_{k}$ is a sequence in $[0,2]$ such that $\sum_{k\in\mathbb{N}}\lambda_{k}(2-\lambda_{k})=+\infty$.
2.2 PnP prior and RED